QUESTION 7 A 2 kg pendulum swings 4 times in 10 seconds. It has a period of QUESTION 8 A 2 kg pendulum swings 4 times in 10 seconds. It has a frequency of cycles per second QUESTION 9 Hz. If a pendulum completes 25 cycles in a minute, its frequency is
Answer #1

(7) time period is time of one period.

so time period(T) = 10/4 = 2.5 sec

(8) frequency is number of cycles per second.

so frequency(f) = 4/10 = 0.4 cycle per second

(9) frequency (f) = 1/T = 1/2.5 = 0.4Hz